# The cosmic trimmer: Black-hole hair in scalar-Gauss-Bonnet gravity is altered by cosmology

### v1

@@ Abstract
Static black holes in general relativity modified by a linear scalar coupling to the Gauss-Bonnet invariant always carry hair. We show that the same mechanism that creates the hair makes it incompatible with a cosmological horizon. Other scalar-tensor models do not have this problem when time-dependence of the scalar provides a natural matching to cosmology. Scalar-Gauss-Bonnet is particularly rigid and such a scenario does not help. An extra operator makes the theory behave like the other models and the cosmological horizon can be accommodated. The hair, however, is drastically altered.
@@ I Introduction
The detection of gravitational waves emitted from the final stages of the inspiral of binary systems of compact objects Abbott et al. (2016, 2017) has opened a new and finally direct avenue for testing the validity of general relativity (GR) as the theory of gravity. These tests, however, require alternative models of gravity which could be used as a laboratory for the prediction of realistic deviations from GR phenomenology. The existence of hair attached to scalar charges carried by compact objects in scalar-tensor theories of gravity provides just such a opportunity. This is the existence of alternative solutions for star-like objects in addition to or instead of the Schwarzschild or Kerr cases possible for the star’s exterior in GR. These hairy (“scalarised”) solutions were discovered in the case of conformally coupled theories of gravity, where it is only possible for neutron stars but not black holes (BH) Bocharova et al. (1970); Damour and Esposito-Farese (1993). Despite the no-hair theorems, see e.g. Bekenstein (1995); Hui and Nicolis (2013), even black holes can carry hair under certain circumstances, for example, when the theory of gravity includes interaction terms between the scalar and higher-order curvature corrections, such as the Gauss-Bonnet term (GB), in principle expected not in the least from effective field theory (EFT) arguments. This alternative phase usually exists only for ranges of BH parameters and appears as a (tachyonic) instability of the GR solution, and can lead to modifications of observable properties of the black holes, such as changes in the spectrum of quasi-normal oscillations or modifications of the BH exterior.
@@
However, the embedding of such setups in cosmological backgrounds presents a challenge. Typically, the parameters which allow for hair in high-curvature environments while permitting a return to GR far from the objects, also cause the equivalent tachyonic instability to appear in the cosmological past. The result is a catastrophic modification of cosmological history compared to the one predicted from GR Babichev et al. (2022). This is incompatible with observations. Perhaps uniquely, the linear coupling to the GB term is not known to suffer from this instability. This theory is quite particular, in that the finiteness of physical observables on the BH horizon in fact requires that BHs carry a charge and the associated hair, making them different than in GR. This has led to a significant body of work studying this model, for a review see Babichev et al. (2016); Lehébel (2018); Herdeiro and Radu (2015) and references within.
@@
Notwithstanding cosmological evolution, attempts have been made to find numerically a static BH solution in scalar-GB gravity in a de Sitter background. They have not been successful Bakopoulos et al. (2019) for so-far unclear reasons.
@@
In this letter, we discuss the case of a linear scalar coupling to GB (sGB) in the context of a black hole embedded within a cosmology under the assumptions of static spacetime and negligible backreaction. We demonstrate that the same finiteness condition that creates the BH hair, makes the cosmological/de Sitter horizon incompatible with it, explaining why such solutions cannot be found. As a result of the shift-symmetry, we can introduce a linear time dependence for the scalar without affecting observables, which has the benefit of being a natural setup in the cosmological embedding. However, this does not solve the problem for this theory. A non-evolving solution in this setup cannot be found.
@@ II Static hair in scalar-Gauss-Bonnet theory
where $\mathcal{G}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2},$ is the Gauss-Bonnet invariant, $\phi$ is dimensionless, and $M_{P}$ is the Planck mass. The corresponding scalar field equation is
$\square\phi+\alpha\mathcal{G}$ $=$ $0\,,$ (2)
The size of the backreaction caused by the sGB term on the background geometry is dictated by the ratio $\alpha/L^{2}$, where $L$ is a characteristic curvature scale of the spacetime Sotiriou and Zhou (2014). Notice that, since sGB is a shift-symmetric Horndeski theory Kobayashi et al. (2011), the scalar field equation (2) can be written as
@@
where the contribution of the sGB term to the current $J^{\mu}$, while not unique in general Yale and Padmanabhan (2011) (see also Creminelli et al. (2020)), it can for example be obtained from the quintic Horndeski current with $G_{5}\sim\log(X)$.
For a static, spherically-symmetric background metric and scalar of the form
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}\,,\qquad\phi=\phi(r)\,,$ (4)
@@
$J_{r}=-\phi^{\prime}+\frac{4\alpha f^{\prime}}{r^{2}f}(1-f)=\frac{C}{r^{2}f}\,,$ (5)
where $C$ is an integration constant to be determined. In order to fix the value of this constant, the standard procedure is to demand the finiteness of $\phi^{\prime}$ and $\phi^{\prime\prime}$ at the black-hole horizon, as proxies for observable quantities. Alternatively, a more covariant and unambiguous way to achieve the same is to require that the kinetic term $X=-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$ remain finite, as it appears for instance in the stress-energy tensor. Near a horizon this goes like
$X=-\frac{f\phi^{\prime 2}}{2}\simeq-\frac{1}{2fr_{h}^{4}}\left(C-4\alpha f^{\prime}_{h}\right)^{2}\,,$ (6)
@@
$C=4\alpha f^{\prime}_{h}\,.$ (7)
Under a test-field approximation, we can take the background metric to be the Schwarzschild one, $f=1-r_{s}/r$, with $r_{s}$ the Schwarzschild radius, and the backreaction being under control as long as $\alpha\ll r_{s}^{2}$. Then it is possible to integrate the scalar field to obtain Sotiriou and Zhou (2014)
$\phi=\frac{4\alpha}{r_{s}r}\left(1+\frac{r_{s}}{2r}+\frac{r_{s}^{2}}{3r^{2}}\right)\,,$ (8)
@@
which asymptotically goes like $C/r$. The black holes in this theory therefore carry a secondary scalar charge, that is, it is not a free parameter but is instead fixed in terms of the black-hole mass $M=M_{P}^{2}r_{s}/2$, as given by Eq. (7). Importantly, it is the presence of a horizon that is responsible for the charge, while horizonless sources such as neutron stars can only support a much-faster-decaying scalar profile Yagi et al. (2016).
@@
Beyond the test-field approximation above, perturbative corrections to the metric have been studied alongside fully self-consistent numerical solutions in Ref. Sotiriou and Zhou (2014), and also considering slow rotation Ayzenberg and Yunes (2014). The observed absence of dephasing on the gravitational-wave waveform during the inspiral phase of binary black-hole mergers has established the most stringent bound on the sGB coupling so far to be $\alpha\lesssim(1.2\,\text{km})^{2}$ Lyu et al. (2022). This justifies the perturbative treatment for $\alpha\ll r_{s}^{2}$ for astrophysical black holes. Further studies have also constrained the validity of the EFT (1), which must have a UV cutoff $\Lambda_{UV}<1/\sqrt{\alpha}$ in order to avoid resolvable superluminal propagation Serra et al. (2022).
@@
Now let us extend the model to include a positive cosmological constant $\Lambda>0$, producing a still static but expanding spacetime with Hubble constant $H$. For observationally allowed values of $\alpha$, $H$ is completely driven by $\Lambda$ and the correction to the speed of gravitational waves is completely negligible. In the test-field approximation, we would expect the Schwarzschild-de Sitter GR solution for the spacetime to carry across, only modifying the metric function to
$f=1-\frac{r_{s}}{r}-H^{2}r^{2}\,.$ (9)
@@
In the astrophysically relevant scenario $Hr_{s}\ll 1$, the Schwarzschild horizon remains at $r_{h}\simeq r_{s}$, but $f$ also has second root and therefore a second horizon – the standard Hubble horizon of cosmology located at $r_{h}\simeq H^{-1}$. For a scalar-field solution valid in the whole observable region, we also have to enforce the finiteness of $X$ on the new cosmological horizon. Eq. (7) then requires that $C$ simultaneously takes different values at each of the horizons, while being a constant and therefore is impossible to satisfy. This is the origin of the non-construction of black-hole solution in the presence of a positive $\Lambda$ in Ref. Bakopoulos et al. (2019). In the next section we will demonstrate that in cosmology the scalar field cannot remain static, and therefore the time dependence must be taken into account when attempting to solve the two-horizon problem.
III Time-dependent hair in cosmologically embedded black-holes
@@
Cosmology suggests a generalisation of this setup which might resolve the problem: it is natural to consider a time-dependent scalar field. We take the simplest ansatz which contains both a time dependence and the spherically symmetric metric required by the local source,
$\phi=q\,t+\varphi(r)\,.$ (10)
@@
$q$ is a new parameter which might allow both the horizons to be compatible with the solution. Owing to the invariance of the action (1) under shifts of $\phi$, such linear time dependence does not give rise to time dependence of any observable quantity. Thus it is compatible with static solutions in principle, as long as the effect of accretion can be neglected. As for the static case, it is again possible to integrate the scalar field equation Eq. (3) once, giving,
$J_{r}=-\varphi^{\prime}+\frac{4\alpha f^{\prime}}{r^{2}f}(1-f)=\frac{C}{r^{2}f}\,,$ (11)
@@
$X=\frac{1}{2}\left(\frac{q^{2}}{f}-f\varphi^{\prime 2}\right)\,,$ (12)
which implies that, unlike in the static case, $\varphi$ can no longer be finite itself if $X(r_{h})=X_{h}$ is to not diverge. Since no observable quantity depends on $\phi$, while $X$ is explicitly contained in at least the energy-momentum tensor, we demand only the finiteness of $X_{h}$ and parameterize the $r$-dependent part of the scalar near any horizon as
$\varphi^{\prime}$ $=$ $\epsilon\frac{q}{f}+\psi^{\prime}(r)\,,$ (13)
@@
$X_{h}=-\epsilon_{h}q\,\psi^{\prime}(r_{h})\,.$ (14)
Then in this near-horizon limit the scalar field equation (11) multiplied by $f$ reads
$\frac{C}{r_{h}^{2}}=J^{r}|_{h}$ $=$ $-\epsilon_{h}q+4\alpha\frac{f^{\prime}_{h}}{r_{h}^{2}}+O(f)\,.$ (15)
Notice that at this order there is no dependence on $X_{h}$, or equivalently $\psi^{\prime}_{h}$. It can now easily be checked that the scalar quantity $J^{2}$ remains finite at the horizons for any value of $C$ [^1] This follows from the fact that, near the horizon, $J_{0}(r_{h})\simeq\pm J^{r}(r_{h})+O(f)$, and therefore $J^{2}\sim 1/q$, which is finite for $q\neq 0$. This is contrary to the static case where the divergence was unavoidable already at the BH horizon Babichev et al. (2016).
@@
$C$ $=$ $r_{s}^{2}\left(-\epsilon_{h}q+\frac{4\alpha}{r_{s}^{3}}\right)\,.$ (16)
Choosing $C$ requires that we pick some $q$ – this is the extra freedom we have introduced. We can do so by examining the cosmological horizon, Eq. (15) naively gives
@@
Seemingly we have found our new solution compatible with both horizons. The question now is how this connects to the cosmological solution.
Let us discuss the theory (1) in a cosmological setting in the absence of a local source. For simplicity we consider de Sitter space,
$ds^{2}=-d\tau^{2}+e^{2H\tau}d{\bf x}^{2}\,,$ (18)
@@
Thus the cosmology imposes that the scalar field always asymptotes to a rolling stationary solution
@@
Translating the cosmological coordinates to the static ones, this implies that in the $r\rightarrow\infty$ limit the cosmological $\phi$ tends to
@@
The only consistent possibility when embedding a BH in cosmology is that the solution sufficiently far away from the BH becomes homogeneous. We thus need to match the scalar hair to this cosmological boundary condition. Comparing this expression for $\varphi^{\prime}$ with the one implied by (11) for $r\rightarrow\infty$ we find that $q$ must take the cosmological value $q_{c}$, while the finiteness condition instead gives
$\frac{q}{q_{c}}=\frac{1}{2Hr_{s}}+1+O(Hr_{s})\gg 1\,.$ (23)
This is vastly different from the one required by cosmology. Moreover, this discrepancy is larger for small black holes. As a matter of principle, the local solutions should not affect the cosmology strongly, but rather the cosmology should represent the aggregated effect of the averaged localized sources – this principle is violated by the static solution in sGB.
This is the main result of this letter: There is an impossibility to construct black-hole solutions in the most simple sGB theory (1) with the proper cosmological asymptotics, when the metric is static and the scalar field time dependent at most in a manner which keeps observables static. The problem results from needing a large charge $C$ to keep the scalar quantities ($X$) finite at the black-hole horizon, which is then incompatible with the same requirement at the cosmological horizon.
One may think that the two-horizon problem could be fixed without much change to the expected physics by breaking the shift-symmetry and adding a potential $V(\phi)$ to the scalar field. In this case the scalar can be asymptotically static at the minimum of $V_{,\phi}+\alpha\mathcal{G}$. Picking $V(\phi)$ such that the mass is of inverse galactic scales would then screen the hair at cosmological distances while giving astrophysics as expected in the static shift-symmetric case. This is true – this modificaton would allow us to use the static solution in the completely static universe. However, during the approach to the asymptotic future de Sitter state, the minimum of the potential evolves and the scalar must roll slowly, giving a small $\dot{\phi}=q$. This would then be enough to put us in the scenario in which the finiteness of $X$ on the horizon would imply that $\phi$ diverges. Since $\phi$ is now a physical observable, this is inconsistent. It remains to be seen whether any such non-shift-symmetric resolution of the problem would not alter the nature of the standard static solution.
Next we investigate an alternative approach to solving the problem that maintains shift-symmetry.

### v2

@@ Abstract
Static black holes in general relativity modified by a linear scalar coupling to the Gauss-Bonnet invariant always carry hair. We show that the same mechanism that creates the hair makes it incompatible with a cosmological horizon. Other scalar-tensor models do not have this problem when time dependence of the scalar provides a natural matching to cosmology. Scalar-Gauss-Bonnet gravity is particularly rigid, and such a scenario does not help. An extra operator makes the theory behave like the other models, and the cosmological horizon can be accommodated. The hair, however, is drastically altered.
@@ I Introduction
The detection of gravitational waves (GWs) emitted from the final stages of the inspiral of binary systems of compact objects [1, 2] has opened a new and finally direct avenue for testing the validity of general relativity (GR) as the theory of gravity. These tests, however, require alternative models of gravity which could be used as a laboratory for the prediction of realistic deviations from GR phenomenology. The existence of hair attached to scalar charges carried by compact objects in scalar-tensor theories of gravity provides just such a opportunity. This is the existence of alternative solutions for starlike objects in addition to or instead of the Schwarzschild or Kerr cases possible for the star’s exterior in GR. These hairy (“scalarized”) solutions were discovered in the case of conformally coupled theories of gravity, where they are only possible for neutron stars but not black holes (BHs) [3, 4] (also see review [5]). Despite the no-hair theorems (see, e.g., [6, 7]), even black holes can carry hair under certain circumstances – for example, when the theory of gravity includes interaction terms between the scalar and higher-order curvature corrections, such as the Gauss-Bonnet term (GB), in principle expected not in the least from effective field theory (EFT) arguments. This alternative phase usually exists only for ranges of BH parameters and appears as a (tachyonic) instability of the GR solution, and it can lead to modifications of observable properties of the black holes, such as changes in the spectrum of quasinormal oscillations or modifications of the BH exterior.
@@
However, the embedding of such setups in cosmological backgrounds presents a challenge. Typically, the parameters which allow for hair in high-curvature environments while permitting a return to GR far from the objects also cause the equivalent tachyonic instability to appear in the early Universe. The result is a catastrophic modification of cosmological history compared to the one predicted from GR [8]. This is incompatible with observations. Modifications of the theory can push the instability into the past, but not remove it [9]. Perhaps uniquely, the linear coupling to the GB term is not known to suffer from this early-Universe instability. This theory is quite particular, in that it is the finiteness of physical observables on the BH horizon rather than the tachyonic instability that requires that BHs carry a charge and the associated hair. This has led to a significant body of work studying this model and its new physics, including the gravitational wave signature [10]. For a review, see [11, 12, 13] and references within.
@@
Notwithstanding cosmological evolution, attempts have been made to find numerically a static BH solution in scalar-GB gravity in a de Sitter background. They have not been successful [14] for so-far unclear reasons.
@@
In this letter, we uncover a new type of inconsistency occurring in the presence of a second horizon, that was not understood previously. We discuss the case of a linear scalar coupling to GB (sGB) in the context of a black hole embedded within a cosmology under the assumptions of static spacetime. This model does not suffer from the tachyonic instability. We nonetheless demonstrate that the same finiteness condition that creates the BH hair makes the cosmological/de Sitter horizon incompatible with it. This two-horizon problem is the reason why such solutions cannot be found. As a result of the shift symmetry, we can introduce a linear time dependence for the scalar without affecting observables, which has the benefit of being a natural setup in the cosmological embedding. However, this does not solve the two-horizon problem for this theory. We conclude that a nonevolving solution in this setup cannot be found.
@@ II Static hair in scalar-Gauss-Bonnet theory
where $\mathcal{G}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ is the Gauss-Bonnet invariant, $\phi$ is dimensionless, and $M_{P}$ is the Planck mass. The corresponding scalar field equation is
$\square\phi+\alpha\mathcal{G}$ $=$ $0\,.$ (2)
The size of the backreaction caused by the sGB term on the background geometry is dictated by the ratio $\alpha/L^{2}$, where $L$ is a characteristic curvature scale of the spacetime [15]. Notice that, since sGB is a shift-symmetric Horndeski theory [16], the scalar field equation (2) can be written as
@@
where the contribution of the sGB term to the current $J^{\mu}$, while not unique in general [17] (see also [18]), it can, for example, be obtained from the quintic Horndeski current with $G_{5}\sim\log(X)$.
For a static, spherically symmetric background metric and scalar of the form
$ds^{2}=-h(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}\,,\qquad\phi=\phi(r)\,,$ (4)
@@
$J_{r}=-\phi^{\prime}+\frac{4\alpha h^{\prime}}{r^{2}h}(1-f)=\frac{C}{r^{2}\sqrt{hf}}\,,$ (5)
where $C$ is an integration constant to be determined. In the static asymptotically flat limit, $f,h\rightarrow 1$, and Eq. (3) implies that $\phi\sim C/r$, with $C$ called the scalar charge of the black hole. In order to fix the value of this constant, the standard procedure is to demand the finiteness of $\phi^{\prime}$ and $\phi^{\prime\prime}$ at the black-hole horizon, as proxies for observable quantities. Alternatively, a more covariant and unambiguous way to achieve the same is to require that the kinetic term $X=-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$ remain finite, as it appears, for instance, in the stress-energy tensor. Near a horizon, this goes like
$X=-\frac{f\phi^{\prime 2}}{2}\simeq-\frac{1}{2hr_{h}^{4}}\left(C-4\alpha\sqrt{\frac{f}{h}}h^{\prime}\right)^{2}\,,$ (6)
@@
$C=4\alpha\sqrt{\frac{f}{h}}h^{\prime}=8\alpha\,\kappa_{\mathrm{BH}}\,,$ (7)
where we have identified the finite quantity $\kappa=\frac{1}{2}\lim_{r\to r_{h}}{\sqrt{\frac{f}{h}}h^{\prime}}$ as the black hole’s surface gravity, which in Schwarzschild is $\kappa_{\text{BH}}=(2r_{s})^{-1}$.
@@
The black holes in this theory therefore carry a secondary scalar charge – that is, it is not a free parameter but is instead fixed in terms of the black hole’s surface gravity, as given by Eq. (7). Importantly, it is the presence of a horizon that is responsible for the charge, while horizonless sources such as neutron stars can only support a much-faster-decaying scalar profile [19].
@@
Perturbative corrections to the metric have been studied, even including slow rotation [20], while fully self-consistent numerical solutions in have been studied in Ref. [15]. The observed absence of dephasing on the gravitational-wave waveform during the inspiral phase of binary black-hole mergers has established the most stringent bound on the sGB coupling so far to be $\alpha\lesssim(1.2\,\text{km})^{2}$ [21]. This justifies a perturbative treatment for $\alpha\ll r_{s}^{2}$ for astrophysical black holes. Further studies have also constrained the validity of the EFT [Eq. (1)], which must have a UV cutoff $\Lambda_{UV}<1/\sqrt{\alpha}$ in order to avoid resolvable superluminal propagation [22].
@@
Now, let us extend the model to include a positive cosmological constant $\Lambda>0$, producing a still static but expanding spacetime with Hubble constant $H$. For observationally allowed values of $\alpha$, $H$ is completely driven by $\Lambda$, and the correction to the speed of gravitational waves is completely negligible. Whenever the backreaction is small, we would expect the Schwarzschild-de Sitter GR solution for the spacetime to be recovered – i.e.,
$h=f=1-\frac{r_{s}}{r}-H^{2}r^{2}\,.$ (8)
@@
In this letter, we only require that the solution (8) be recovered asymptotically at cosmological distances, but we allow for backreaction close to the black hole. In the astrophysically relevant scenario $Hr_{s}\ll 1$, the black-hole horizon remains at $r_{h}\simeq r_{s}$, but $f,h$ also have second roots and therefore a second horizon – the standard Hubble horizon of cosmology located at $r_{h}\simeq H^{-1}$. For a scalar-field solution valid in the whole observable region, we also have to enforce the finiteness of $X$ on the new cosmological horizon. Equation (7) then requires that $C$ simultaneously take different values at each of the horizons, while being a constant, and therefore it is impossible to satisfy. This is the origin of the nonconstruction of the black-hole solution in the presence of a positive $\Lambda$ in Ref. [14]. In the next section, we will demonstrate that in cosmology, the scalar field cannot remain static, and therefore the time dependence must be taken into account when attempting to solve the two-horizon problem.
III Time-dependent hair in cosmologically embedded black holes
@@
Cosmology suggests a generalization of this setup which might resolve the problem: it is natural to consider a time-dependent scalar field. We take the simplest ansatz which contains both a time dependence and the spherically symmetric metric required by the local source,
$\phi=q\,t+\varphi(r)\,,$ (9)
@@
where $q$ is a new parameter which might allow both the horizons to be compatible with the solution. Owing to the invariance of the action (1) under shifts of $\phi$, such linear time dependence does not give rise to time dependence of any observable quantity. Thus, it is compatible with static solutions in principle, as long as the effect of accretion can be neglected. As for the static case, it is again possible to integrate the scalar field equation Eq. (3) once, giving
$J_{r}=-\varphi^{\prime}+\frac{4\alpha h^{\prime}}{r^{2}h}(1-f)=\frac{C}{r^{2}\sqrt{hf}}\,,$ (10)
@@
$X=\frac{1}{2}\left(\frac{q^{2}}{h}-f\varphi^{\prime 2}\right)\,,$ (11)
which implies that, unlike in the static case, $\varphi$ can no longer be finite itself if $X(r_{h})=X_{h}$ is not to diverge. Since no observable quantity depends on $\phi$, while $X$ is explicitly contained in at least the energy-momentum tensor, we demand only the finiteness of $X_{h}$ and parametrize the $r$-dependent part of the scalar near any horizon as
$\varphi^{\prime}$ $=$ $\epsilon\frac{q}{\sqrt{hf}}+\psi^{\prime}(r)\,,$ (12)
@@
$X_{h}=-\epsilon_{h}q\sqrt{\frac{f}{h}}\,\psi^{\prime}(r_{h})\,.$ (13)
Then in this near-horizon limit the scalar field equation (10) multiplied by $\sqrt{hf}$ reads
$\frac{C}{r_{h}^{2}}=\sqrt{\frac{h}{f}}J^{r}|_{h}$ $=$ $-\epsilon_{h}q+4\alpha\sqrt{\frac{f}{h}}\frac{h^{\prime}}{r_{h}^{2}}+O(h,f)\,.$ (14)
Notice that at this order, there is no dependence on $X_{h}$, or equivalently $\psi^{\prime}_{h}$. It can now easily be checked that the scalar quantity $J^{2}$ remains finite at the horizons for any value of $C$ [^1] This follows, provided that $\frac{f^{\prime}}{f}-\frac{h^{\prime}}{h}$ is regular, which is in turn necessary to avoid a curvature singularity. Then $J^{2}\sim 1/q$, which is finite for $q\neq 0$. This is contrary to the static case, where the divergence was unavoidable already at the BH horizon [11].
@@
$C$ $=$ $r_{s}^{2}\left(-\epsilon_{h}q+\frac{8\alpha}{r_{s}^{2}}\kappa_{\mathrm{BH}}\right)\,.$ (15)
Choosing $C$ requires that we pick some $q$ – this is the extra freedom we have introduced. We can do so by examining the cosmological horizon: Eq. (14) naively gives
@@
where it is assumed that the geometry approaches de Sitter. Seemingly, we have found our new solution compatible with both horizons. The question now is how this connects to the cosmological solution.
Let us discuss the theory [Eq. (1)] in a cosmological setting in the absence of a local source. For simplicity, we consider de Sitter space,
$ds^{2}=-d\tau^{2}+e^{2H\tau}\left(d\rho^{2}+\rho^{2}d\Omega^{2}\right)\,,$ (17)
@@
Thus, the cosmology imposes that the scalar field always asymptotes to a rolling stationary solution
@@
Translating the cosmological coordinates to the static ones, this implies that in the $r\rightarrow\infty$ limit, the cosmological $\phi$ tends toward
@@
The only consistent possibility when embedding a BH in cosmology is that the solution sufficiently far away from the BH becomes homogeneous. We thus need to match the scalar hair to this cosmological boundary condition [Eq. (20)] sufficiently far away. This requires an $r$ dependence as in Eq. (21), which is satisfied by the solution of Eq. (10) for $r\rightarrow\infty$, while the “local” time dependence $q$ only contributes decaying terms in observable quantities, [^2] We thank Dražen Glavan for useful comments about this point. e.g., $X-X_{c}\simeq e^{-2H\tau}(q-q_{c})^{2}/2(H\rho)^{2}$. A similar behavior was observed for Brans-Dicke theory [23]. However, in the case of sGB, the finiteness condition fixes $q$ as
$\frac{q}{q_{c}}=\frac{1}{2(Hr_{s})}\frac{\kappa_{\mathrm{BH}}}{\kappa_{s}}+1+O(Hr_{s})\gg 1.$ (22)
This is vastly different from the cosmological value $q_{c}$, implying either (a) a very slow approach to homogeneity only at distances $H\rho\sim 1/(Hr_{s})\gg 1$, or (b) a backreaction so strong as to completely screen the gravitational effect of the black hole – i.e., $\kappa_{\textrm{BH}}\lesssim(Hr_{s})\kappa_{s}\ll\kappa_{s}$. Moreover, this tension is larger for small black holes. As a matter of principle, the local solutions should not affect the cosmology strongly, but rather the cosmology should represent the aggregated effect of the averaged localized sources – this principle is violated by the stationary solution in sGB.
This is the main result of this letter: There is an impossibility to construct black-hole solutions in the most simple sGB theory [Eq. (1)] with a physically acceptable approach to the homogeneous cosmological asymptotics, when the metric is static and the scalar field is time dependent at most in a manner which keeps observables static. The problem stems from the $r_{s}^{-1}$ scaling of the required values of $C$ and $q$ to keep scalar quantities ($X$) finite at both the black-hole and cosmological horizons, making it incompatible with having both the usual notion of a black hole and cosmological homogeneity at reasonable separation.
One may think that the two-horizon problem could be fixed without much change to the expected physics by breaking the shift symmetry and adding a potential $V(\phi)$ to the scalar field. In this case, the scalar can be asymptotically static at the minimum of $V_{,\phi}+\alpha\mathcal{G}$. Picking $V(\phi)$ such that the mass is of inverse galactic scales would then screen the hair at cosmological distances while giving astrophysics as expected in the static shift-symmetric case. This is true – this modification would allow us to use the static solution in the completely static universe. However, during the approach to the asymptotic future de Sitter state, the minimum of the potential evolves and the scalar must roll slowly, giving a small $\dot{\phi}=q$. This would then be enough to put us in the scenario in which the finiteness of $X$ on the horizon would imply that $\phi$ diverges. Since $\phi$ is now a physical observable, this is inconsistent. It remains to be seen whether any such non-shift-symmetric resolution of the problem would not alter the nature of the standard static solution. A similar conclusion can be reached for a mass term provided by a coupling to the Ricci scalar $\phi^{2}R$.
Next, we investigate an alternative approach to solving the problem that maintains shift symmetry.